OpenStudy (anonymous):

How would I rewrite y= square root of (4x=16)- 3?

OpenStudy (anonymous):

\[y = \sqrt{4x=16} -3?\]

OpenStudy (anonymous):

Oh and describing the graph too, I know I have to do the factor tree?

OpenStudy (anonymous):

Is that what you have?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Check again, it doesn't make sense

OpenStudy (anonymous):

\[y=\sqrt{4x+16}-13\]

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\large y = a\sqrt{x-h} + k\) is a transformation for its parent : \(\large y = \sqrt{x}\)

ganeshie8 (ganeshie8):

change ur equation to above form

ganeshie8 (ganeshie8):

\(\large y=\sqrt{4x+16}-13\) \(\large y=\sqrt{4(x+4)}-13\) \(\large y=2\sqrt{x- (-4)}-13\)

OpenStudy (anonymous):

I don't know what my Y is

ganeshie8 (ganeshie8):

^^thats the form they want u change the given equation to

ganeshie8 (ganeshie8):

next, graph it

ganeshie8 (ganeshie8):

-4 moves the parent graph to left by 4 units -13 moves the parent graph down by 13 units

ganeshie8 (ganeshie8):

the factor, 2 in front of radical stretches the graph vertically by a factor of 2